train a and b are traveling in the same direction on parallel tracks. Train A is traveling at 60 miles per hr and train B is traveling at 80 miles per hr. Train A passes at 2:15 Pm and train b passes t 2:27 Pm. What time will train B catch up to A
Answer by mananth(16949)


Train A 60 mph
Train B = 80 mph
02:15
02:27
Difference in time= 00:12
Train A will have covered 12 miles before Train B has started
catch up distance= 12 miles
catch up speed = 80 -60 mph
catch up speed = 20 mph
Catchup time = catchup distance/catch up speed
catch up time= 12 / 20
catch up time= 0.6 hours
catch up time= 36 minutes
02:27+36= = 03:03 pm